A solution contains 2P.5 g of cholecalciferol which is also known as Vitamin D3 (C27H44O) in 4500 mL of methanol (CH3OH). [ Here P is the last three digits of your Student ID) [Density of methanol at 25 C is 0.792 g/mL]

(a) Calculate the molarity of this solution

(b) Calculate the mole fraction of Vit D3 and methanol

(c) Is above mixture a dilute or concentrated solution? Explain your understanding within 60 words


1
Expert's answer
2021-05-03T07:40:23-0400

Assuming the last three digits of your Student ID is 011


(a) number of moles = 2011.5/384.6 = 5.23 moles

molarity = number of moles/volume = 5.23/4500 × 1000 = 1.162M


(b) mass of methanol = 0.792 × 4500 = 3564g

number of moles of methanol = 3654/32 = 111.375 moles


molar fraction of Vit D3 = 5.23/116.605 = 0.045

molar fraction of methanol = 111.375/116.695 = 0.972


(c) concentrated solution because of its molarity.